OpenStudy (anonymous):

Simplify:a+4(6a+2b^2+(a^2-6b^2) : )

OpenStudy (anonymous):

=a+4 (6a + 2b^2 + a ^2 - 6b^2) =a+4( a^2 + 6a - 4b^2) =a( a^2 + 6a - 4b^2) + 4( a^2 + 6a - 4b^2) =a^3 + 6a^2 -4b^2a + 4a^2 + 24a - 16b^2 =a^3 + 10a^2 +24a - 4b^2a - 16b^2

OpenStudy (anonymous):

Thats the answer ??

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

. i dont think so : /

OpenStudy (anonymous):

then it should be: 25a - 16b^2 +4a^2

OpenStudy (anonymous):

Not that either . the answer choices dont have that answer . . .

OpenStudy (anonymous):

http://www.webmath.com/polymult.html type your problem here

OpenStudy (anonymous):

. im at school. its a blocked site . thats what it says here .

OpenStudy (anonymous):

oh okay